# Characteristic quasi-polynomials of deletions of Shi arrangements of type B and their period collapse

### v2

@@ Abstract
Characteristic quasi-polynomials are the enumerative functions counting the number of elements in the complement of hyperplane arrangements modulo positive integers. In this paper, we compute the characteristic quasi-polynomials of the restriction of the Shi arrangement of type B by one given hyperplane. As a corollary, we determine whether period collapse occurs in the characteristic quasi-polynomial of the deletion of the Shi arrangement of type B. This implies the solution for the conjecture posed by Higashitani, Tran and Yoshinaga in the case of the Shi arrangement of type B.
@@ 1 Introduction
In the theory of hyperplane arrangements, the characteristic polynomial of a hyperplane arrangement $\mathcal{A}$ plays an important role and reflects many algebraic/combinatorial properties on $\mathcal{A}$. In [3] and [4], Kamiya, Takemura and Terao introduce a new invariant, which can be regarded as an extension of the characteristic polynomial. It is proved in [3] and [4] that the number of points contained in the complement of $\mathcal{A}$ whose elements are regarded as the points in $(\mathbb{Z}/q\mathbb{Z})^{m}$ becomes a quasi-polynomial. This quasi-polynomial is called a characteristic quasi-polynomial. (The detailed definition will be explained in Section 2.1.) Since then, characteristic quasi-polynomials have been investigated for several classes of hyperplane arrangements. For example, the characteristic polynomials of the extended Linial arrangements are used to prove the conjecture by Postnikov and Stanley [7] (see [9] and [8]). Moreover, characteristic quasi-polynomials of the extended Shi arrangements are also computed as follows:
@@
Now, we introduce the main object of this paper. Recall that the Shi arrangement $\mathcal{B}_{m}$ of type B is defined by
@@
where $m]=\{1,2,\ldots,m\}$.
On the other hand, we say that a period collapse occurs if the minimum period of $\mathcal{A}$ becomes strictly less than the lcm period of $\mathcal{A}$. (See Section 2.2 for more details.) It is showed in [2] that for any positive integers $m$, $p$ and $s$ such that $s$ divides $p$, there exists a non-central hyperplane $m$-arrangement whose characteristic quasi-polynomial has lcm period $p$ but has minimum period $s$. Roughly speaking, this means that “any possible period collapse” can occur. Here, Theorem 1.1 says that the extended Shi arrangement has the characteristic quasi-polynomial as just a polynomial, i.e., minimum period becomes $1$. On the other hand, the lcm period of the extended Shi arrangement of type B is known to be $2$. Hence, period collapse occurs in the extended Shi arrangement of type B.

### v3

@@ Abstract
Characteristic quasi-polynomials are the enumerative functions counting the number of elements in the complement of hyperplane arrangements modulo positive integers. A notable phenomenon in this context is period collapse, where the quasi-polynomial reduces to a polynomial or has a smaller period than the lcm period. In this paper, we compute the characteristic quasi-polynomials of the restriction of the Shi arrangement of type B by one given hyperplane. As a corollary, we completely determine whether period collapse occurs in the characteristic quasi-polynomial of the deletion of the Shi arrangement of type B. This implies the solution for the conjecture posed by Higashitani, Tran and Yoshinaga in this case.
@@ 1 Introduction
Quasi-polynomials arise naturally in various areas of mathematics, including Ehrhart theory, representation theory, and the study of arithmetic and combinatorial invariants associated with discrete geometric structures. One of the most remarkable phenomena in this context is period collapse, where a quasi-polynomial exhibits a strictly smaller period than expected, or even reduces to an ordinary polynomial. Despite its apparent simplicity, period collapse remains poorly understood in general, and only a limited number of explicit and structurally rich examples have been identified so far. In this paper, we focus on the characteristic quasi-polynomial of hyperplane arrangements and analyze the occurrence of period collapse, offering new contributions to this context of research.
In the theory of hyperplane arrangements, the characteristic polynomial of an arrangement $\mathcal{A}$ is a fundamental invariant encoding deep algebraic and combinatorial information. In [3, 4], Kamiya, Takemura, and Terao introduced a natural generalization of the characteristic polynomial, called the characteristic quasi-polynomial. More precisely, they showed that for a hyperplane arrangement $\mathcal{A}$ defined over the integers, the number of points in the complement of $\mathcal{A}$ when regarded as a subset of $(\mathbb{Z}/q\mathbb{Z})^{m}$ is a quasi-polynomial of $q$. (The precise definition will be recalled in Section 2.1.) This invariant provides a bridge between hyperplane arrangements and arithmetic aspects of quasi-polynomial behavior. In fact, recent work by Liu, Tran, and Yoshinaga [5] shows that characteristic quasi-polynomials are closely related to invariants of associated toric arrangements, further clarifying the combinatorial and arithmetic information captured by these invariants.
Since its introduction, the characteristic quasi-polynomial has been computed and studied for several important classes of hyperplane arrangements. For instance, characteristic quasi-polynomials of extended Linial arrangements play a key role in the proof of a conjecture of Postnikov and Stanley [7] (see [9] and [8]). Characteristic quasi-polynomials of extended Shi arrangements have also been investigated, revealing rich arithmetic and combinatorial structures and providing concrete instances where periodic phenomena can be analyzed explicitly:
@@
Recall that the Shi arrangement $\mathcal{B}_{m}$ of type B is defined by
@@
where $m]=\{1,2,\ldots,m\}$. In this paper, we study deletions of Shi arrangements associated with the root system of type $B$. These arrangements form a natural and tractable family within the broader framework of hyperplane arrangements, where explicit computations of characteristic quasi-polynomials are possible. Our main results provide closed formulas for these quasi-polynomials and show that, in this setting, period collapse occurs systematically. Thus, the Shi arrangements of type $B$ offer a new and computable source of examples illuminating the general phenomenon of period collapse in characteristic quasi-polynomials.
To be precise, we say that a period collapse occurs for a characteristic quasi-polynomial of a hyperplane arrangement $\mathcal{A}$ if its minimum period is strictly smaller than the lcm period. (See Section 2.2 for precise definitions and further details.) It is shown in [2] that for any positive integers $m$, $p$ and $s$ such that $s$ divides $p$, there exists a non-central hyperplane $m$-arrangement whose characteristic quasi-polynomial has lcm period $p$ but has minimum period $s$. Roughly speaking, this means that “any possible period collapse” can occur. In addition, Theorem 1.1 shows that the characteristic quasi-polynomial of the extended Shi arrangement is in fact a polynomial, i.e., its minimum period equals $1$. On the other hand, the lcm period of the extended Shi arrangement of type B is known to be $2$. Hence, period collapse occurs in the extended Shi arrangement of type B.